Use the distributive property to remove the parentheses in the following expression. Then simplify your result if possible.
1/4 (20+16)

Answer by fractalier(2101)

1/4*(20 + 16) =
(1/4)(20) + (1/4)(16) =
20/4 + 16/4 =
5 + 4 =
9